### v1

# MetaAligner: Conditional Weak-to-Strong Correction for Generalizable Multi-Objective Alignment of Language Models

Abstract
Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses. MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 11 policy models with up to 63$\times$ more parameters, and outperforms previous alignment methods with down to 22.27$\times$ less computational resources. The model also accurately aligns with unseen objectives, marking the first step towards generalizable multi-objective preference alignment. Models and codes will be available at: https://github.com/SteveKGYang/MetaAligner
1 Introduction
The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020), where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model (Ouyang et al., 2022).
Algorithm Paradigm Reward-Free Align. Multi-Objective Align. Policy-Agnostic Align. Generalizability
RLHF (Ouyang et al., 2022) PPO
MORLHF (Li et al., 2020) PPO
MODPO (Zhou et al., 2023) DPO
CPO (Guo et al., 2024) SFT, DPO
RiC (Yang et al., 2024) SFT
Aligner (Ji et al., 2024) SFT
MetaAligner SFT
Table 1: A checklist for key characteristics of previous preference alignment methods and MetaAligner. "Align." denotes "Alignment".
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences (Bai et al., 2022; Rame et al., 2024). These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) (Sener and Koltun, 2018; Li et al., 2020; Rame et al., 2024). Due to its substantial computational cost (Li et al., 2020; Rame et al., 2024) and the unstable nature of the PPO (Schulman et al., 2017; Kumar et al., 2020; Rafailov et al., 2024), other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) (Zhou et al., 2023) or supervised fine-tuning (SFT)-based methods (Yang et al., 2024; Guo et al., 2024), which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023; Chiang et al., 2023; Team et al., 2024). (2) all target models are statically aligned on pre-determined (e.g. Helpfulness, Harmlessness, Honesty (Zhou et al., 2023; Guo et al., 2024)) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access. This is achieved by extending the Aligner (Ji et al., 2024) to multi-objective alignment scenarios, which are not directly solvable by Aligner itself. Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives. The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction. That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering. During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections. As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment. The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning (Kojima et al., 2022). The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. A checklist of MetaAligner’s key advantages over previous methods is presented in Table 1.
Through this work, we make three main contributions:
  • We propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or even accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods with down to 22.27× less computational resources and higher stability.
  • We explore utilizing MetaAligner to exert zero-shot preference alignment for unseen objectives. To the best of our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner performs effective alignment for 3 unseen objectives while maintaining performance on aligned objectives.
  • We examine MetaAligner on the IMHI benchmark (Yang et al., 2023b), which consists of 9 sub-tasks for mental health analysis and 20K testing samples. Experimental results show that MetaAligner trained on merely 7.2K human preference data improves win-rate on multiple objectives across 11 policy models with sizes ranging from 2B to 70B, substantially enhancing responses from policy models with up to 63$\times$ more parameters.
Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF (Li et al., 2020), (B) SFT-based methods (Yang et al., 2024; Guo et al., 2024), and (C) Meta-Objective Aligner. SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
2 Multi-Objective Alignment
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplay such as contradiction (e.g. helpfulness and harmlessness) and dependence (e.g. correctness and informativeness). Multi-objective preference alignment tackles this challenge by optimizing on multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$. $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry (Bradley and Terry, 1952) model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\frac{1}{1+exp(-(r_{i}(q,y_{1})-r_{i}(q,y_{2})))}$ (1)
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards (Sener and Koltun, 2018; Li et al., 2020) with a KL-divergence regularization (see Figure 1 (A)):
$\small\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector. In another paradigm, Yang et al. (2024); Guo et al. (2024) directly build alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q,\mathbf{R}(q,y))\right]$ (3)
where objectives and their corresponding reward values are described with text markers and combined into queries with a static prompting template. Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable.
3 Meta-Objective Aligner
This section introduces the proposed methodology for MetaAligner, which follows a three-stage paradigm: (1) Dynamic multi-objective dataset construction; (2) a three-step model training process; (3) policy-agnostic and generalizable inference for alignment. The paradigm is illustrated in Figure 1 (C).
3.1 Dynamic Multi-Objective Dataset
We propose constructing a dynamic multi-objective dataset to trigger MetaAligner’s ability for reward-free and flexible adjustment of alignment objectives. Specifically, a typical multi-objective preference alignment dataset with $N$ objectives $\mathcal{D}_{n}$ is formalized as $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$, where $R_{i}=[r_{i1},...,r_{iN}]^{\mathrm{T}}$ and $r_{ij}\in\{\succ,\prec,=\}$. We define a text marker for each objective: $<R>:[<r_{1}>,...,<r_{N}>]$, where $<r_{j}>$ denotes a natural language description marking the $i$-th objective. Six examples of such text markers are presented in Figure 1 (C). With a pre-defined prompting template $\mathcal{T}(q,y,\mathcal{O},t)$, we build a preference subset $\mathcal{D}_{p}$ and another equal-preference subset $\mathcal{D}_{e}$ from $\mathcal{D}_{n}$, where $\mathcal{D}_{p}$ includes all contrastive response pairs and $\mathcal{D}_{e}$ includes all equal-labelled response pairs. We utilize the following prompting template for building the datasets:
[$\mathcal{T}(q,y,\mathcal{O},t)$] Edit the following Question-Answer pair to make it {$t$} considering the following objectives $\mathcal{O}$} | Question: {$q$} | Answer: {$y$} | Edit:
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text markers for the target objectives, and $t\in\{equal,better\}$ depends on the building subset. The detailed construction process is described in Algorithm 1.
Algorithm 1 Training data construction
1:Raw datasets: $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$; Objective text markers: $[<r_{1}>,...,<r_{N}>]$; Prompting template: $\mathcal{T}(q,y,\mathcal{O},t)$
2:Preference subset $\mathcal{D}_{p}$; All-equal subset $\mathcal{D}_{e}$.
3:$\mathcal{D}_{p}\leftarrow\varnothing$, $\mathcal{D}_{e}\leftarrow\varnothing$triangleright Initialize the 2 subsets.
4:for $i\in\{1,...,n\}$ do
5:    $\mathcal{O}_{\succ}\leftarrow\varnothing$, $\mathcal{O}_{\prec}\leftarrow\varnothing$, $\mathcal{O}_{=}\leftarrow\varnothing$
6:    for $j\in\{1,...,N\}$ do
7:         if $r_{ij}$ is $\succ$ $\triangleright$ Preferred obj. for $y_{1}$.
8:             $\mathcal{O}_{\succ}\leftarrow\mathcal{O}_{\succ}\cup\{<r_{j}>\}$
9:         else if $r_{ij}$ is $\prec$ $\triangleright$ Preferred obj. for $y_{2}$.
10:             $\mathcal{O}_{\prec}\leftarrow\mathcal{O}_{\prec}\cup\{<r_{j}>\}$
11:         $\triangleright$ Equally preferred objectives.
12:             $\mathcal{O}_{=}\leftarrow\mathcal{O}_{=}\cup\{<r_{j}>\}$
13:         end if
14:    end for
15:    if $\mathcal{O}_{\succ}\neq\varnothing$ $\triangleright$ Build $y_{1}$-preferred data.
16:         $t\leftarrow<better>$
17:         $\mathcal{O}_{\succ}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\succ})$
18:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\succ},t),y_{i1})\}$
19:    end if
20:    if $\mathcal{O}_{\prec}\neq\varnothing$ $\triangleright$ Build $y_{2}$-preferred data.
21:         $t\leftarrow<better>$
22:         $\mathcal{O}_{\prec}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\prec})$
23:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i1},\mathcal{O}_{\prec},t),y_{i2})\}$
24:    end if
25:    if $\mathcal{O}_{=}\neq\varnothing$ $\triangleright$ Build equal data.
26:         $t\leftarrow<equal>$
27:         $\mathcal{O}_{=}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{=})$
28:         $\mathcal{D}_{e}\leftarrow\mathcal{D}_{e}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{=},t),y_{i1})\}$
29:    end if
30:end for
Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping (Yang et al., 2024) process in previous SFT-based multi-objective alignment methods.
3.2 MetaAligner Derivation
Based on the dynamic multi-objective dataset, we achieve multi-objective alignment by introducing MetaAligner, a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the initial outputs as follows:
$\pi^{*}(y|q)=\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\pi_{\phi}(y_{0}|q)$ (4)
where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$. We achieve alignment by directly employing SFT on the target response. We have the standard cross-entropy loss as the training objective:
$\mathop{argmin}\limits_{\theta,\phi}$ $\mathcal{L}(\theta,\phi;\mathcal{D})=-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\pi^{*}(y|q)\right]$ (5)
=-= - $\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-$
$\mathbb{E}_{q\sim\mathcal{D}}\left[log\pi_{\phi}(y_{0}|q)\right]$
We fuse the parameters of the policy model, thus excluding $\phi$ from the weight update process. As only $\theta$ derivatives are calculated during back-propagation, we eliminate the second term in Eqn 5 and simplify the training objective as:
$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as output from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs (Achiam et al., 2023; Team et al., 2023).
3.3 Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a three-step paradigm based on the objective function in Eqn. 6:
Step 1: warm-up.
The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies (He et al., 2016; Ji et al., 2024) to familiarize the model with identity mapping and facilitate convergence. Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
Step 2: equal-preference modelling.
Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, in objectives such as Correctness and Honesty (Yang et al., 2023b; Li et al., 2023b), equal preferences are common and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
Step 3: preference alignment
This stage fine-tunes the MetaAligner on the preference subset $\mathcal{D}_{p}$, which instructs the model to perform conditional weak-to-strong correction on the dis-preferred response and achieve alignment to the preferred response on the specified objectives.
3.4 Generalizable Inference
During inference, MetaAligner achieves alignment following the sampling process as in Eqn. 4, where unaligned outputs, sampled from the target policy model, are used as the input for conditional weak-to-strong correction. With the prompting-based paradigm, the target objectives for MetaAligner also become expandable and generalizable, a key advantage over previous alignment methods (Zhou et al., 2023; Yang et al., 2024; Guo et al., 2024). The generalizability is two-fold: Firstly, users can manipulate the target objectives by adjusting combinations of text markers in the objective set $\mathcal{O}$. For example, in alignment with objectives 1 and 3, we can incorporate the corresponding text markers as follows:
$\mathcal{O}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:<r_{1}>}$ (7)
Secondly, the prompt-based objectives statement enables flexible adjustment of text markers for existing objectives and injections of unseen objectives. Following the last example, we have two unseen alignment objectives $r_{1}^{*}:<r_{1}^{*}>$, $r_{2}^{*}:<r_{2}^{*}>$, and an updated text marker $<r_{1}^{*}>$ for aligned objective 1. We can perform zero-shot alignment on these objectives by adjusting $\mathcal{O}$ as follows:
$\mathcal{O}^{*}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:}{\color[rgb]{1,0,0}<r_{1}^{*}>;r_{1}^{*}:<r_{1}^{*}>;r_{2}^{*}:<r_{2}^{*}>}$ (8)
This simple pattern can theoretically lead to unlimited simultaneous alignment objectives. We expect MetaAligner to make generalizable weak-to-strong corrections under these unseen conditions via its in-context learning ability. This advancement marks a new exploration into generalizable multi-objective preference alignment.
MetaAligner Policy Model CAMS DR Dreaddit IRF loneliness MultiWD SAD SWMH T-SID Overall
1.1B LLaMA2-Chat-7B -3.4% +28.0% +12.67% +45.67% +38.0% +34.67% +23.33% +33.0% +31.33% +27.04%
LLaMA2-Chat-13B -25.0% +27.67% +22.67% +52.33% +32.0% +31.33% +23.33% +23.67% +36.67% +29.7%
LLaMA2-Chat-70B +5.33% +35.0% +19.0% +46.33% +42.0% -3.0% +8.0% +0.33% +3.0% +21.18%
Gemma-instruct-2B +32.0% +4.33% +37.0% +40.0% +34.0% +61.0% +45.0% +52.0% +27.33% +38.77%
Gemma-instruct-7B +16.33% +51.33% +42.67% +44.67% +51.0% +55.33% +49.33% +40.0% +53.67% +44.92%
MentaLLaMA-7B +13.33% +22.67% +23.33% +47.33% +39.67% +39.33% +33.33% +30.67% +41.0% +32.29%
MentaLLaMA-13B +21.34% +31.0% +39.0% +29.67% +42.33% +47.66% +20.67% +34.33% +33.0% +36.03%
MentaLLaMA-33B -5.44% -0.33% -5.33% +30.67% +3.0% +14.0% +24.33% +7.0% +2.33% +6.66%
Vicuna-7B +6.0% -9.33% +24.0% +70.67% +17.67% +33.33% +17.33% +22.0% +42.0% +24.85%
Vicuna-13B +6.0% -9.0% +13.67% +71.67% +20.33% +32.67% +21.0% +2.0% +16.67% +24.97%
Vicuna-33B +29.33% +2.0% -1.0% +41.0% +10.0% +9.67% +1.33% -9.67% +1.33% +9.33%
7B LLaMA2-Chat-7B +7.67% +47.0% +13.0% +33.0% +38.33% +31.0% +17.0% +38.67% +49.33% +30.55%
LLaMA2-Chat-13B -12.34% +51.67% +18.0% +37.67% +43.33% +29.33% +22.33% +18.0% +38.34% +27.37%
LLaMA2-Chat-70B +15.66% +40.53% +9.0% +28.33% +25.34% +6.67% +31.33% +10.33% +10.66% +20.67%
Gemma-instruct-2B +45.0% +8.33% +30.0% +57.0% +39.0% +48.66% +57.0% +52.0% +36.33% +41.11%
Gemma-instruct-7B +32.0% +42.67% +20.67% +45.0% +36.33% +43.0% +46.0% +40.67% +20.7% +38.0%
MentaLLaMA-7B +20.66% +50.67% +19.66% +35.66% +30.0% +27.0% +40.33% +36.34% +41.0% +35.48%
MentaLLaMA-13B +25.0% +45.66% +43.66% +34.34% +36.33% +30.33% +48.0% +42.33% +36.0% +37.97%
MentaLLaMA-33B -5.33% +2.66% -9.66% +2.34% +21.33% -4.67% +20.0% +8.67% +3.33% +4.22%
Vicuna-7B +22.33% +7.67% +4.0% +40.0% +20.0% +14.0% +6.0% +7.67% +15.0% +15.19%
Vicuna-13B +1.0% +48.0% +8.0% +67.0% +33.0% +12.0% +15.0% +23.0% +22.0% +32.33%
Vicuna-33B -2.0% +54.0% +10.0% +62.0% +37.0% +29.0% +10.0% +15.0% +12.0% +25.22%
13B LLaMA2-Chat-7B +27.33% +45.0% +23.33% +62.67% +54.33% +59.34% +52.0% +55.33% +70.33% +49.96%
LLaMA2-Chat-13B +5.34% +47.0% +22.67% +65.33% +56.33% +52.33% +31.33% +32.0% +56.67% +35.55%
LLaMA2-Chat-70B +28.0% +59.0% +23.33% +57.33% +53.0% +7.0% +27.33% +23.0% +23.67% +31.18%
Gemma-instruct-2B +52.66% +19.66% +41.33% +60.67% +50.67% +79.0% +53.67% +63.33% +45% +51.74%
Gemma-instruct-7B +35.33% +52.0% +39.0% +55.34% +50.0% +64.0% +56.33% +49.33% +61.0% +50.36%
MentaLLaMA-7B +40.67% +38.34% +29.33% +65.67% +38.34% +57.0% +46.66% +49.0% +61.67% +47.4%
MentaLLaMA-13B +34.0% +27.0% +39.0% +54.33% +35.33% +49.33% +41.67% +46.67% +45.66% +43.62%
MentaLLaMA-33B +2.67% -8.0% +2.67% +30.67% -8.0% +23.67% +29.33% +28.67% +20.0% +11.2%
Vicuna-7B +23.33% +17.67% +8.0% +63.67% +35.67% +36.33% +12.33% +18.0% +37.0% +28.07%
Vicuna-13B +15.0% -7.33% +2.0% +67.67% +24.33% +40.67% +19.0% +4.0% +17.67% +20.29%
Vicuna-33B -6.0% +54.0% +36.0% +79.0% +38.0% +29.0% +24.0% +21.0% +24.0% +33.22%
Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models. The GPT-4 judge considers 3 objectives: Correctness, Informativeness, and Professionality. The figures show the advantage of aligned outputs over the policy model outputs on win rate. Best values for each MetaAligner model are highlighted in bold.
4 Experiments
4.1 Experimental Settings
Datasets
We train and evaluate the performance of MetaAligner on the IMHI dataset (Yang et al., 2023b). For the training data, we invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives with the following text markers: (1) Correctness: the response should make correct predictions in the corresponding sub-tasks; (2) informativeness: the response should express clear logic and provide consistent supporting evidence; (3) Professionality: the generated responses should provide supportive evidence with high quality and reliability. For evaluation, we perform alignment on the IMHI benchmark, which consists of 9 sub-tasks with 20K query-response pairs.
Models
We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B (Zhang et al., 2024) and LLaMA2-Chat-(7B, 13B) (Touvron et al., 2023) foundation models. We utilize MetaAligner to perform multi-objective alignment on the following policy models: LLaMA2-Chat-(7B,13B,70B) (Touvron et al., 2023), Gemma-instruct-(2B,7B) (Team et al., 2024), MentaLLaMA-(7B, 13B, 33B) (Yang et al., 2023b), and Vicuna-(7B, 13B, 33B) (Chiang et al., 2023). Among the policy models, MentaLLaMA models are instruction-tuned on mental health-related data, and other models are tuned on general-domain instruction data.
Evaluation Metric
We quantify the alignment performance of model outputs by comparing their win rate against the ground-truth response provided by the IMHI benchmark, on the corresponding objectives. Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works (Guo et al., 2024; Sun et al., 2023; Li et al., 2023a), to perform the judgments, where each target response, query, ground-truth response, and objectives are provided via prompt engineering. GPT-4 is required to select the response with higher alignment with the specified objectives in the prompt, or indicate a tied performance.
Figure 2: Objective-wise win-rate comparisons based on different policy models. MetaAligner-7B with different levels of alignment are tested on Correctness, Informativeness, and Professionality, compared with 3 different policy models. The win rates are obtained via the judgments of GPT-4 on the corresponding objectives.
4.2 Overall Performance
We stack MetaAligner on different policy models to perform alignment on all 3 objectives: Correctness, Informativeness, and Professionality. The overall performance on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 2.
According to the results, the MetaAligner models achieve substantial improvement in overall performance on all 11 policy models, with an average of 26.89% advantage on win rates for MetaAligner-1.1B, 28.01% for MetaAligner-7B, and 36.6% for MetaAligner-13B. These results show the general effectiveness of one MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment. MetaAligner also greatly improves performance on each sub-task. For example, MetaAligner-7B outperforms the unaligned outputs by over 25% on 7 sub-tasks. These results indicate that MetaAligner alignment can be effectively adapted to tasks that require different knowledge and response formats.
From the policy model scale perspective, MetaAligner provides successful alignments to models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself. In the extreme case, MetaAligner-1.1B advances the win-rate of LLaMA2-Chat-70B outputs by 21.18%, a policy model with 63$\times$ more parameters. These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow. Besides the general-domain foundation models, MetaAligner also improves the performance by an average of 28.32% on MentaLLaMA models, which are fine-tuned on mental health analysis tasks. These results show that MetaAligner can make reasonable corrections on weak responses while maintaining their expertise from domain-specific policy models.
Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase. These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner. Fortunately, MetaAligner’s capabilities also show scalability. Increasing the size of its base model leads to a higher win-rate advantage on most policy models. Specifically, MetaAligner-7B outperforms or achieves comparable performance to MetaAligner-1.1B on 9 policy models, and MetaAligner-13B surpasses MetaAligner-7B on 10 policy models by an average of 12.58%. These observations motivate further explorations in model scale-performance balance for MetaAligner.
Figure 3: Generalizable alignment on 3 unseen objectives based on LLaMA2-Chat-70B outputs. In the x-axis, "+" denotes the addition of a new objective into the objective set $\mathcal{O}$. "★" denotes the unaligned win rates for LLaMA2-Chat-70B outputs. "-." lines identify win rate fluctuations before alignment on the corresponding objective, and solid lines identify fluctuations after alignment.
4.3 Objective-Wise Evaluation
We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives. We utilize MetaAligner to perform 5 levels of alignments: No alignment, aligning on each objective (Correctness, Informativeness, and Professionality), and full alignment. The objective-wise win rates are shown in Figure 2. We have the following observations:
Objective-wise alignment significantly improves the target objective and benefits the performance on other objectives. For example, aligning on Informativeness with MetaAligner achieves the best performance on Informativeness, and increases win rates on Correctness and Professionality over the unaligned outputs. This tendency holds with all policy models and other objectives. These results further prove the complex interplay within multi-objective alignment, where positive correlations also exist, in addition to contradictions (Guo et al., 2024).
Aligning on all objectives provides balanced performance. According to the results, full alignment displays competitive performance on all 3 objectives and generally outperforms unaligned outputs and aligned outputs from other objectives, even comparable to aligned outputs from the same objective in some cases. The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, enabling the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
MetaAligner performs comprehensive alignment on different sub-tasks. With most policy models, MetaAligner significantly outperforms the unaligned responses on all 9 sub-tasks, regardless of the alignment objectives. These results indicate the general effectiveness of our dynamic-objective method on tasks demanding various expertise (Yang et al., 2023b). However, MetaAligner still struggles to align certain sub-tasks, especially with strong policy models. For example, MetaAligner outputs fail to outperform MentaLLaMA-13B on the CAMS sub-task in all objectives. These results show enhancements in domain knowledge, with techniques such as retrieval-augment generation (Lewis et al., 2020), a promising future direction for improvement.
Algorithm Space Complexity Correctness Informativeness Professionality Overall
MORLHF 22.27$\times$ -6.31% +33.7% +16.31% +15.08%
MODPO 17.27$\times$ +26.31% +30.1% +33.12% +32.9%
SFT 11.82$\times$ +19.22% +24.89% +25.6% +26.62
MetaAligner-1.1B 1$\times$ +26.11% +35.67% +13.11% +29.7%
MetaAligner-7B 6.36$\times$ +28.89% +24.92% +29.07% +27.37%
MetaAligner-13B 11.82$\times$ +31.33% +45.0% +43.66% +35.55%
Table 3: Win rates of previous alignment methods and MetaAligner models based on the LLaMA2-Chat-13B policy model. Space complexity is calculated based on the memory consumption of training MetaAligner-1.1B.
4.4 Generalizable Alignment
In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 3 less relevant objectives that are unseen during its training phases: (1) Completeness (Yang et al., 2023a): the response should cover all relevant aspects of the original query; (2) Honesty (Guo et al., 2024): the response should speak in truth that is consistent with the query and commonsense. (3) Safety (Sun et al., 2023): The response should avoid content that is offensive, discriminatory, or harmful.
2,700 random-selected responses from the IMHI benchmark are re-aligned with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one. Their performance is presented in Figure 3. According to the results, we have the following conclusions:
MetaAligner performs effective zero-shot alignment for unseen objectives. With all MetaAligner models, incorporating an unseen objective significantly improves its corresponding win rate. For example, MetaAligner-7B improves win rate by 40.7% on Completeness, 28.89% on Honesty, and 17.56% on Safety compared to each of these objectives not introduced. These results prove the viability of generalizable alignment with the in-context learning ability of MetaAligner. In addition, alignment performance on supervised objectives (Correctness, Informativeness, Professionality) generally surpasses unseen objectives on all tested models, showing that MetaAligner also learns objective-specific features during training.
Performance on aligned objectives is maintained with more unseen alignment objectives. As each objective is aligned, its win rate surges, then stabilizes as long as the objective is included in $\mathcal{O}$. These results prove MetaAligner to perform reliable alignment with the expansion of objectives. However, aligning on new objectives can affect other unaligned objectives. For example, aligning on Correctness with MetaAligner-1.1B benefits its win rate by 11.22%, but harms performance on 3 other unaligned objectives. These problems are due to the limited ability of smaller MetaAligner models and the controversial dependence of certain multi-objective alignments (Guo et al., 2024), where we observe an alleviation as the model size increases.
MetaAligner’s generalizability also shows scalability. Performance on the 3 unseen objectives increases with the scale-up of MetaAligner model size, where MetaAligner-1.1B improves 12.7% on average, MetaAligner-7B improves 29.05%, and MetaAligner-13B improves 36.69%. This scalability is attributed to the growing in-context learning ability of larger base models, which motivates further explorations into the correlation between generalizable alignment and model scales.
4.5 MetaAligner vs. MORLHF/MODPO/SFT
We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on 2700 random-selected responses from the IMHI benchmark. We implement the linear scalarization method for MORLHF. We follow previous settings of realizing MODPO (Guo et al., 2024) by pre-pending the objective markers. Due to the closed-source nature of current SFT-based methods, we only replicate stage 1 where details are clear, and utilize the preference-to-reward mapping method proposed in Yang et al. (2024). We select LLaMA2-Chat-13B as the policy model, and the results are presented in Table 3.
According to the results, MetaAligner-13B achieves over 30% of win rates on all objectives, and generally outperforms all previous methods by 10.68% on average. This advantage is achieved with equal or less space complexity than previous algorithms, showing the high efficiency of the proposed method. As the base model size further reduces, MetaAligner-1.1B and MetaAligner-7B show lower, but still competitive performance to the 13B model. Impressively, MetaAligner-1.1B outperforms both MORLHF and SFT-based methods by 8.85% on average, and approaches MODPO with merely a 3.2% gap, but with 11.82$\times$-22.27$\times$ less computational resources. This huge advantage indicates a high prospect for applying MetaAligner in low-resource scenarios. For previous methods, MODPO and SFT can provide stable alignment for multiple objectives, but with high space complexity. MORLHF bears the highest resource consumption, but achieves the lowest 15.08% overall win rate. Its training process is also unstable, including a 6.31% decrease in the Correctness objective.
5 Related Work
5.1 Large Language Models
Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language Achiam et al. (2023); Team et al. (2023); Touvron et al. (2023); Luo et al. (2023). Notable examples include commercially available LLMs like ChatGPT, GPT-4 Achiam et al. (2023), and Gemini Team et al. (2023), as well as leading-edge open-source models like LLaMA2 Touvron et al. (2023) and Mistral Jiang et al. (2023). Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information Zhang et al. (2023); Xie et al. (2023), flattery, and deception, raising concerns about their potential negative impacts on society and associated risks Bommasani et al. (2021). To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning Ji et al. (2023).
5.2 Preference Alignment
Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty. These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022). Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences. This limitation has motivated the exploration of multi-objective RLHF (MORLHF) Sener and Koltun (2018); Li et al. (2020); Rame et al. (2024), by linear scalarizations of multiple rewards Sener and Koltun (2018); Li et al. (2020) and interpolations of LLM weights trained from diverse reward models Rame et al. (2024). However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) Zhou et al. (2023) without the RL paradigm. Zhou et al. (2023) proposed multi-objective direct preference optimization (MODPO), which extended DPO algorithm (Rafailov et al., 2024) to combine multiple objectives with specific weightings. Another paradigm is the SFT-based methods Yang et al. (2024); Guo et al. (2024), which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO. These methods also facilitated objective-wise controllable generation during inference.
6 Discussion
Conclusion.
This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment. Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives. Thorough investigations on LLMs ranging from 2B to 70B proved MetaAligner’s overall effectiveness in multi-objective alignment for various policy models and objective-wise alignment. and Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align more objectives.
Limitation.
Though achieving outstanding performance, MetaAligner still bears limitations. Firstly, similar to the vanilla Aligner module Ji et al. (2024), stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference, which affects actual model deployment, especially for scenarios such as local deployment on mobile devices. Secondly, MetaAligner still provides limited improvement for large-scale domain-specific policy models, such as MentaLLaMA-33B, in domain knowledge-intense tasks such as mental health analysis. These results reflect the importance of expertise knowledge for multi-objective alignment, where MetaAligner still requires further enhancements. Thirdly, due to limited resources, we only tested the generalizability of MetaAligner on 3 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives.
Future Work.
Firstly, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation (Lewis et al., 2020). Secondly, we will dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance. We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 6 objectives). This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.