### v2

# MetaAligner: Towards Generalizable Multi-Objective Alignment of Language Models

Abstract
Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses. MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and outperforms previous alignment methods with down to 15.71$\times$ less GPU training hours. The model also effectively aligns unseen objectives, marking the first step towards generalizable multi-objective preference alignment. This project is open-sourced at: https://github.com/SteveKGYang/MetaAligner
1 Introduction
The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [39, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model [23].
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences [4, 25]. These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) [27, 20, 25]. Due to its substantial computational cost [20, 25] and the unstable nature of the PPO [26, 17, 24], other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [38, 11] or supervised fine-tuning (SFT)-based methods [35, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g. "Helpful", "Harmless", "Honest" [38, 11]) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
Algorithm Paradigm Reward-Free Align. Multi-Objective Align. Policy-Agnostic Align. Generalizability
RLHF [23] PPO
MORLHF [20] PPO
MODPO [11, 38] SFT, DPO
RiC [35] SFT
Aligner [13] SFT
MetaAligner SFT
Table 1: A checklist for key characteristics of previous preference alignment methods and MetaAligner. "Align." denotes "Alignment".
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access [13]. Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives. The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction. That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering. During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections. As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment. The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning [16]. The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. A checklist of MetaAligner’s advantages over previous methods is presented in Table 1.
Through this work, we make three main contributions: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods with down to 15.71× less GPU training hours and higher stability; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives. To our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner can simultaneously perform effective alignment for 6 unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on 3 preference alignment datasets. Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5 [22] and Claude-3 [3].
Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF [20], (B) SFT-based methods [35, 11], and (C) Meta-Objective Aligner. SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
2 Multi-Objective Alignment of Language Models
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplays such as contradiction (e.g. "Helpful" and "Harmless") and dependence (e.g. "Correct" and "Informative"). Multi-objective preference alignment tackles this challenge by optimizing multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$. $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\frac{1}{1+exp(-(r_{i}(q,y_{1})-r_{i}(q,y_{2})))}$ (1)
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization (see Figure 1 (A)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector. Another paradigm [35, 11] directly built alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q,\mathbf{R}(q,y))\right]$ (3)
where objectives and their corresponding reward values are described with text markers and combined into queries with a static prompting template. Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable.
3 Meta-Objective Aligner
This section introduces the proposed methodology for MetaAligner, which follows a 3-stage paradigm: (1) dynamic multi-objective dataset construction; (2) a 3-step model training process; (3) generalizable inference for alignment. The paradigm is illustrated in Figure 1 (C).
3.1 Dynamic Multi-Objective Dataset
[…]
[$\mathcal{T}(q,y,\mathcal{O},t)$] Edit the following Question-Answer pair to make it {$t$} considering the following objectives $\mathcal{O}$} | Question: {$q$} | Answer: {$y$} | Edit:
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the building subset. The detailed construction process is described in Algorithm 1. A full list of the utilized prompting templates is presented in Appendix B.
Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping [35] process in previous SFT-based multi-objective alignment methods.
3.2 MetaAligner Derivation
Based on the dynamic multi-objective dataset, we achieve multi-objective alignment by introducing MetaAligner, a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the initial outputs as follows:
$\pi^{*}(y|q)=\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\pi_{\phi}(y_{0}|q)$ (4)
where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$. We achieve alignment by directly employing SFT on the target response. We have the standard cross-entropy loss as the training objective:
$\mathop{argmin}\limits_{\theta,\phi}$ $\mathcal{L}(\theta,\phi;\mathcal{D})=-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\pi^{*}(y|q)\right]$ (5)
=-= - $\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-$
$\mathbb{E}_{q\sim\mathcal{D}}\left[log\pi_{\phi}(y_{0}|q)\right]$
We fuse the parameters of the policy model, thus excluding $\phi$ from the weight update process. As only $\theta$ derivatives are calculated during back-propagation, we eliminate the second term in Eqn 5 and simplify the training objective as:
$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as samples from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs [2, 22, 3].
3.3 Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a 3-step paradigm based on the objective function in Eqn. 6:
Step 1: warm-up.
The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies [12, 13] to familiarize the model with identity mapping and facilitate convergence. Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
Step 2: equal-preference modeling.
Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, equal preferences are common in many scenarios [34, 8] and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
Step 3: preference alignment
This stage fine-tunes the MetaAligner on the preference subset $\mathcal{D}_{p}$, which instructs the model to perform conditional weak-to-strong correction on the dis-preferred response and achieve alignment to the preferred response on the specified objectives.
3.4 Generalizable Inference
[…]
This simple pattern can theoretically lead to unlimited simultaneous alignment objectives. We expect MetaAligner to make generalizable weak-to-strong corrections under these unseen conditions via its in-context learning ability. This advancement marks a new exploration into generalizable multi-objective preference alignment.
4 Experiments
4.1 Experimental Settings
Datasets.
We train and evaluate on 3 alignment datasets: (1) HH-RLHF [4]: a large-scale dataset with 160K prompts and corresponding response pairs. We follow previous works [35] and utilize 3 open-sourced reward models on "Harmless", "Helpful", and "Humor" to assess and rank the responses in their corresponding objectives; (2) UltraFeedback [8]: a multi-aspect alignment dataset with 64K prompts with preferences obtained from GPT-4. We select "Instruction following", "Honest", "Truthful", and "Helpful" objectives during training; (3) IMHI: we evaluate MetaAligner on domain knowledge-intense scenarios by creating another alignment dataset based on the IMHI dataset [34] targeting interpretable mental health analysis, which consists of 9 mental health analysis sub-tasks. We invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives: "Correct", "Informative", and "Professional".
Models.
We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B [36] and LLaMA2-(7B, 13B) [31] foundation models. We utilize MetaAligner to perform multi-objective alignment on the following open-source policy models: LLaMA2-Chat-(7B,13B,70B) [31], Gemma-instruct-(2B,7B) [30], and Vicuna-(7B, 13B, 33B) [7]. We also align 2 advanced close-sourced foundation models: GPT-3.5 [22] and Claude-3-Sonnet [3], where model parameters are inaccessible.
Evaluation Metric.
On each objective, we quantify the alignment performance of model outputs by comparing their win rates against the ground-truth response provided by the benchmark datasets. Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works [11, 29, 19], to perform the judgments. Each target response, ground-truth response, query, and evaluated objectives are provided via prompt engineering. GPT-4 is required to compare and select the response with higher alignment to the specified objectives or indicate a tied performance.
More details about the training process, model cards, dataset statistics, IMHI dataset annotation, and evaluation settings are presented in Appendix C.
HH-RLHF UltraFeedback IMHI
MetaAligner Policy Model Harmless Helpful Humor IF Honest Truthful Helpful Correct Informative Professional
1.1B LLaMA2-Chat-7B +10.0% +20.0% +14.75% +11.0% +15.0% +14.33% +9.0% +18.33% +20.55% +31.67%
LLaMA2-Chat-13B +10.75% +9.08% +13.25% +8.66% +15.34% +16.33% +7.67% +11.11% +8.33% +25.0%
LLaMA2-Chat-70B +6.58% +7.42% +22.58% +6.0% +12.67% +17.33% +16.33% +8.33% +14.23% +17.23%
Gemma-instruct-2B +8.5% +12.25% +12.33% +14.67% +14.67% +13.0% +5.33% 15.55% +35.55% +37.23%
Gemma-instruct-7B +4.0% +7.75% +23.17% +9.0% +10.0% +4.67% +14.0% +18.9% +31.12% +36.11%
Vicuna-7B +11.5% +10.83% +20.33% +11.33% +13.33% +12.33% +7.0% +10.0% +7.22% +6.33%
Vicuna-13B +7.42% +13.0% +19.17% +11.66% +14.34% +15.33% +10.0% +12.22% +7.78% +3.34%
Vicuna-33B +8.5% +2.59% +23.83% +8.0% +11.67% +6.33% +6.67% +8.34% +4.44% +6.12%
GPT-3.5 +1.42% +7.5% +17.84% +5.0% +5.0% +3.66% +1.0% +9.67% +1.33% +9.33%
Claude-3-Sonnet -3.83% +1.58% +13.17% +4.67% +2.67% +2.67% +3.0% +7.0% +2.33% +6.66%
7B LLaMA2-Chat-7B +25.0% +27.0% +20.75% +34.66% +36.0% +37.0% +28.0% +21.67% +32.22% +43.89%
LLaMA2-Chat-13B +28.75% +20.58% +18.25% 34.0% +37.34% +37.66% +23.3% +25.56% +30.0% +33.89%
LLaMA2-Chat-70B +16.58% +14.42% +29.08% +31.0% +27.0% +31.33% +17.0% +20.56% +17.23% +21.67%
Gemma-instruct-2B +20.0% +18.75% +17.83% +41.33% +40.67% +42.33% +31.33% +25.0% +50.55% +51.67%
Gemma-instruct-7B +11.0% +23.25% +26.67% +33.67% +35.34% +31.0% +29.0% +35.01% +52.23% +56.11%
Vicuna-7B +19.5% +18.83% +27.33% +38.0% +39.0% +37.0% +32.33% +23.33% +22.78% +23.33%
Vicuna-13B +14.92% +21.0% +30.67% +34.66% +40.0% +39.67% +36.34% +25.55% +20.0% +15.01%
Vicuna-33B +28.0% +17.09% +30.83% +30.0% +37.34% +32.33% +29.33% +11.11% +16.11% +8.34%
GPT-3.5 +15.92% +21.5% +22.84% +29.99% +30.34% +28.0% +14.34% +18.67% +16.33% +14.22%
Claude-3-Sonnet +19.17% +19.08% +26.17% +22.33% +21.0% +21.67% +19.0% +11.33% +19.33% +11.33%
13B LLaMA2-Chat-7B +21.0% +30.5% +23.75% +51.83% +47.5% +45.33% +38.67% +28.33% +38.33% +50.56%
LLaMA2-Chat-13B +22.75% +16.08% +12.25% +46.33% +48.67% +46.83% +41.17% +30.56% +37.22% +40.56%
LLaMA2-Chat-70B +16.58% +11.92% +23.08% +44.33% +35.0% +45.5% +24.0% +31.67% +30.56% +36.12%
Gemma-instruct-2B +27.5% +20.75% +24.33% +55.0% +44.67% +51.33% +36.83% +35.55% +63.33% +65.0%
Gemma-instruct-7B +14.0% +27.75% +25.17% +42.0% +40.17% +35.17% +31.17% +34.45% +50.0% +49.44%
Vicuna-7B +18.0% +12.83% +30.33% +41.5% +39.83% +44.33% +37.5% +24.44% +23.33% +21.11%
Vicuna-13B +17.92% +29.0% +32.67% +47.33% +49.17% +47.0% +40.67% +28.33% +23.34% +18.9%
Vicuna-33B +31.5% +17.09% +27.83% +50.5% +53.17% +45.83% +38.5% +23.89% +23.89% +14.45%
GPT-3.5 +19.92% +25.0% +23.84% +40.33% +40.17% +36.83% +23.67% +26.67% +25.66% +33.62%
Claude-3-Sonnet +27.17% +23.58% +27.17% +38.5% +39.5% +37.67% +29.83% +28.67% +20.0% +11.2%
Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on 3 datasets over different policy models. The responses are simultaneously aligned on all objectives, then tested on each objective. "IF" denotes the "Instruction following" objective. "+" shows the advantage of aligned outputs over the unaligned outputs on win rates against the golden responses.
4.2 Overall Performance
MetaAligner-(1.1B, 7B, 13B) performance on 3 alignment datasets are shown in Table 2. According to the results, the MetaAligner models achieve substantial improvement for most objectives and policy models. For example, on UltraFeedback, there is an average of 11.47% advantage for MetaAligner-1.1B on "Honest", 34.39% for MetaAligner-7B, and 43.79% for MetaAligner-13B. These results show the general effectiveness of MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment. On the mental health analysis benchmark IMHI, MetaAligner models also show remarkable win rates on all objectives, proving their effectiveness in performing multi-objective alignment in domain knowledge-intense scenarios. We further evaluate MetaAligner on each IMHI sub-task and the results are shown in Appendix D.
From the policy model scale perspective, MetaAligner provides successful alignments to open-sourced models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself. In the extreme case, MetaAligner-1.1B advances the win rates of LLaMA2-Chat-70B outputs, a policy model with 63$\times$ more parameters, by an average of 12.19% on HH-RLHF, 13.08% on UltraFedback, and 13.26% on IMHI. These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow. MetaAligner also significantly improves performance on API-based models: GPT-3.5 and Claude-3-Sonnet. These results prove its potential for application in close-sourced scenarios and effective multi-objective alignment of state-of-the-art policy models.
Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase. These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner. Fortunately, MetaAligner’s capabilities also show scalability. Increasing the size of its base model leads to a higher win-rate advantage on most policy models. For example, on UltraFeedback, MetaAligner-7B outperforms MetaAligner-1.1B on all 10 policy models, and MetaAligner-13B further surpasses MetaAligner-7B by an average of 12.58%. These observations motivate further explorations in model scale-performance balance for MetaAligner.
4.3 Evaluations of Objective-Wise Alignment
Figure 2: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives. The results are the performance of MetaAligner-7B on LLaMA2-Chat-70B outputs from the UltraFeedback test set.
We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives. We utilize MetaAligner to perform 6 levels of alignments: unaligned, aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and full alignment. We leverage GPT-4 to score the responses ranging from 0 to 10. Experimental details and more results are shown in Appendix E. The results are shown in Figure 2. We have the following observations:
Objective-wise alignment improves performance on the target objective and benefits the performance on other objectives. For example, Aligning on "Instruction following" achieves the best GPT4 score distribution on the "Instruction following" evaluation results. It also significantly increases GPT4 scores on "Helpful", "Honest", and "Truthful" over the unaligned responses. This tendency holds with other policy models and alignment objectives. These results further prove the complex interplay within multi-objective alignment, where positive correlations and contradictions [11] co-exist.
Full alignment on all objectives provides balanced performance. According to the results, full alignment displays competitive performance on all 4 objectives. Generally, it outperforms unaligned outputs and aligned outputs from other objectives, even comparable to those from the same objective, such as in "Honest". The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, training the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
Figure 3: Zero-shot alignment on 6 unseen objectives based on LLaMA2-Chat-70B outputs of UltraFeedback. In the x-axis, "Aligned Obj." denotes the 4 supervised objectives ($\diamond$" markers), and "+" denotes further addition of an unseen objective ($\circ$" markers). $\star$" denotes the unaligned win rates for the unseen objectives, "-." lines identify win rate fluctuations before alignment, and solid lines identify fluctuations after alignment.
4.4 Generalizable Alignment
In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 6 less relevant objectives that are unseen during its training phases: Specific, Factual, Readable, Fair, Repeat, and Length [10]. More details about these objectives are in Appendix B.
We randomly select 2,700 queries from the UltraFeedback dataset and re-align the LLaMA2-Chat-70B outputs with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one, with 10 aligned objectives in total. Their win rates on each objective over the golden responses are presented in Figure 3. According to the results, we have the following conclusions:
MetaAligner performs effective zero-shot alignment for unseen objectives. With most MetaAligner models, incorporating an unseen objective into the objective set significantly improves its corresponding win rate. For example, MetaAligner-7B improves by 25.17% on "Specific", 14.5% on "Factual", and 17.5% on "Readable" compared to each of these objectives unaligned. These results prove the viability of generalizable alignment with the in-context learning ability. However, the win rates on supervised objectives ("Instruction following", "Helpful", "Honest", and "Truthful") generally surpass unseen objectives, showing that supervised learning remains more effective in multi-objective preference alignment compared to in-context learning.
Performance on aligned objectives is maintained with more unseen alignment objectives. As each objective is aligned, its win rate surges, stabilizing as long as it is included. On simultaneously aligning 10 objectives, MetaAligner-7B outperforms LLaMA2-Chat-70B outputs by an average of 14.25% on unseen objectives. These results prove MetaAligner to perform overall reliable alignment with the expansion of objectives. However, aligning on new objectives can affect other objectives. For example, aligning on "Fair" with MetaAligner-(7B, 13B) benefits its win rates, but harms performance on objectives such as "Readable" and "Factual". The decreases are due to the "alignment tax" [11], where enhancements in one objective can diminish performance in other objectives due to their controversial nature.
MetaAligner’s generalizability shows scalability. Performance on the 6 unseen objectives increases with the scale-up of MetaAligner model size. MetaAligner-1.1B provides limited alignment on most unseen objectives, but MetaAligner-7B extends the win rates to an average of 48.5%, and MetaAligner-13B further reaches 61.25%. MetaAligner-13B also more effectively aligns objectives such as "Length", where smaller models perform badly. This scalability is attributed to larger foundation models’ growing in-context learning ability, which enables accurate interpretations of the objective descriptions and instructions. These observations motivate further explorations into the correlation between generalizable alignment and model scales in future work.
HH-RLHF UltraFeedback
Algorithm GPU Hours Harmless Helpful Humour Avg. IF Honest Truthful Helpful Avg.
MORLHF 1892.3 62.83% 51.2% 77.5% 63.84% 32.18% 33.7% 26.1% 33.7% 31.42%
MODPO 405.9 65.0% 64.0% 78.0% 69.0% 30.82% 43.4% 37.19% 25.0% 34.1%
SFT 247.34 66.5% 75.0% 76.5% 72.67% 27.0% 36.5% 26.0% 36.5% 31.5%
MetaAligner-1.1B 120.48 62.5% 75.0% 77.0% 71.5% 27.67% 27.0% 33.0% 25.33% 28.25%
MetaAligner-7B 242.68 77.5% 82.0% 83.0% 80.83% 51.33% 48.0% 55.67% 44.33% 49.83%
MetaAligner-13B 403.44 73.5% 85.5% 86.0% 81.67% 68.5% 59.5% 64.0% 55.0% 61.75%
Table 3: Comparisons of win rates between alignment methods based on the LLaMA2-Chat-7B policy model. "GPU hours" records the summed GPU running time on all datasets.
4.5 MetaAligner vs. MORLHF/MODPO/SFT-based Method
We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on the HH-RLHF and UltraFeedback datasets, where explicit rewards are available. We implement the linear scalarization method for MORLHF, the CDPO [11] realization of MODPO, and RiC [35] realization of the SFT-based method. We select LLaMA2-Chat-7B as the policy model, and the results are presented in Table 3. Appendix F presents details about the baseline model implementations and GPU hours calculations.
According to the results, MetaAligner-13B significantly outperforms all other methods with an average of 81.67% win rate on HH-RLHF and 61.75% on UltraFeedback, showing the general advantage of the conditional weak-to-strong correction paradigm. As the base model size reduces, MetaAligner shows decreased but still competitive performance compared to other baseline models, but achieved with less memory consumption and GPU training hours. Impressively, the MetaAligner-1.1B model achieves comparable average performance to MORLHF, MODPO, and SFT-based methods on both datasets, but with 2.05$\times$-15.67$\times$ less GPU training hours and 6.36$\times$ smaller size than the LLaMA2-Chat-7B policy model. These facts indicate the high efficiency of MetaAligner algorithms and a prospect for application in low-resource scenarios. Compared to previous methods, MetaAligner models can also achieve balanced performances in objective-wise evaluations, including contradictory objectives such as "Harmless" and "Helpful", without requiring explicit hyper-parameter tuning for achieving Pareto optimal solutions [11, 35, 20]. In other methods, inappropriate heuristic preference weight selection can lead to serious performance degradation in certain objectives. For example, with a uniform distribution of preference weights, the performance of MORLHF on "Helpful" falls to 51.2%, a huge gap to other methods.
5 Related Work
5.1 Large Language Models
Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language [2, 31, 21]. Notable examples include commercially available LLMs like ChatGPT, GPT-4 [2], and Claude-3 [3], as well as leading-edge open-source models like LLaMA2 [31] and Mistral [15]. Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information [37, 32], flattery, and deception, raising concerns about their potential negative impacts on society and associated risks [5]. To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning [14].
5.2 Human Preference Alignment
Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty. These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) [23]. Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences. This limitation has motivated the exploration of multi-objective RLHF (MORLHF) [27, 20, 25], by linear scalarizations of multiple rewards [27, 20] and interpolations of LLM weights trained from diverse reward models [25]. However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) [38] without the RL paradigm. Multi-objective direct preference optimization (MODPO)[38] extended DPO algorithm [24] to combine multiple objectives with specific weightings. Another paradigm is the SFT-based methods [35, 11], which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO. These methods also facilitated objective-wise controllable generation during inference.
6 Discussions
Conclusion.
This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment. Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives. Thorough investigations on various LLM policy models proved MetaAligner’s overall effectiveness in multi-objective alignment and objective-wise alignment. Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align multiple objectives.
Limitation and Future Work.
Firstly, stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference [13], which affects model deployment, especially for scenarios such as local deployment on mobile devices. Secondly, due to limited resources, we only tested the generalizability of MetaAligner on 6 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives. In future work, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation [18]. We will also dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance. We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 10 objectives). This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.
Appendix A Ethics and Impacts
A.1 Licenses
We leveraged 3 publicly available datasets to build our dynamic multi-objective datasets: HH-RLHF, UltraFeedback, and IMHI. The licenses of the datasets and the 3 publicly available reward models we used to annotate the HH-RLHF dataset are available in Table 5. In Sec. 4.5, we implement the reward assignment scripts for HH-RLHF and MORLHF based on the released codes of Yang et al. [35], which is available at Github. The MODPO and MORLHF codes are also based on the OpenRLHF framework under the Apache-2.0 license. The code, data, and the MetaAligner models will be released for replication of the results and future usage, under the MIT license.
A.2 Broader Impacts
In this work, MetaAligner provides an effective and model-agnostic method for generalizable and expandable alignment of LLM outputs with multiple human expectations. It has great potential to develop AI assistants more accurately aligned with human intentions and social values. However, the prompt-based nature of the objective selection process facilitates the customization of new alignment objectives, which can be easily misused to align responses with malicious objectives (e.g. sexism, racism, suicide ideation) via adjusting the objective descriptions and utilizing the in-context learning ability of MetaAligner. These actions can lead to harmful outputs from MetaAligner. As the authors of MetaAligner, we are dedicated to developing safe and fair AI technology to benefit the common welfare of our society. We condemn any malicious use of MetaAligner and advocate for its responsible and ethical applications. In addition, as MetaAligner performs alignment in a plug-and-play manner on top of the policy models, deployment of this technology can increase the overall inference cost of AI assistants and carbon emissions. These disadvantages can affect the long-term goals of developing green AI systems and equitable access to AI to benefit all of humanity.
A.3 Safeguards
This released codes, data, and MetaAligner models are provided for research only. None of the material constitutes actual diagnosis or advice, and help-seekers should get assistance from professional psychiatrists or clinical practitioners. No warranties, express or implied, are offered regarding the accuracy, completeness, or utility of the responses and explanations. The authors and contributors are not responsible for any errors, omissions, or any consequences arising from the use of the information herein. Users should exercise their own judgment and consult professionals before making any clinical-related decisions. The use of the software and information contained in this paper is entirely at the user’s own risk.
The collected queries to build our IMHI preference dataset are from the publicly available IMHI dataset [34], and we strictly follow the privacy protocols and ethical principles to protect user privacy and guarantee that anonymity is properly applied in all the mental health-related texts. In addition, to minimize misuse, all examples provided in our paper are paraphrased and obfuscated utilizing the moderate disguising scheme.
In addition, recent studies have indicated LLMs may introduce some potential bias, such as gender gaps. Meanwhile, some incorrect prediction results, inappropriate explanations, and over-generalization also illustrate the potential risks of current LLMs. Therefore, there are still many challenges in applying the models to real scenarios.
By using or accessing the information in this paper, the users agree to indemnify, defend, and hold harmless the authors, contributors, and any affiliated organizations or persons from any and all claims or damages.
Appendix B Text Descriptions and Prompting Templates
The text descriptions for all tested objectives in this paper are included in Table 4. The descriptions are determined via the definition of Wikipedia and further polished to fit the tasks. During alignment, the text descriptions are combined with the text marker of the objectives to provide clear instructions on the target. The aligned objectives are obtained from the annotations of each alignment dataset, and the unaligned objectives are selected from the definitions of previous works [10].
On building the dynamic multi-objective dataset, we carefully define prompting templates to trigger the conditional weak-to-strong correction ability of MetaAligner. A full list of the used templates is shown below. Specifically, for the preference subset of the HH-RLHF dataset [4], since the model is required to improve the response considering a multi-turn dialogue history, we have:
[$\mathcal{T}(q,y,\mathcal{O},"better")$] You are an assistant to human. You will be provided with a context and an answer. Consider the context, then edit the answer to improve it considering these aspects: $\{\mathcal{O}\}$ | Context: {q} | Answer: {y} | Edit:
For the equal-preference subset, we have:
[$\mathcal{T}(q,y,\mathcal{O},"equal")$] You are an assistant to human. You will be provided with a context and an answer. Consider the context, then edit the answer to make it equal considering these aspects: $\mathcal{O}$} | Context: {q} | Answer: {y} | Edit:
In the UltraFeedback dataset [8], the model is required to improve the response considering a single query. For the preference subset, we have:
[$\mathcal{T}(q,y,\mathcal{O},"better")$] You are an assistant to human. You will be provided with a query and an answer. Consider the query, then edit the answer to improve it considering these aspects: $\mathcal{O}$} | Query: {q} | Answer: {y} | Edit:
For the equal-preference subset, we have:
[$\mathcal{T}(q,y,\mathcal{O},"equal")$] You are an assistant to human. You will be provided with a query and an answer. Consider the query, then edit the answer to make it equal considering these aspects: $\mathcal{O}$} | Query: {q} | Answer: {y} | Edit:
Objectives $r$ Text Description $r\right>$
Aligned Objectives
Harmless The response should avoid content that is offensive, discriminatory, or harmful.
Helpful The response should provide useful resources and suggestions to the user.
Humor The response should be cheerful and amusing.
Instruction following The response should carefully follow the instructions of the query.
Honest The response should not tell lies
Truthful The response should actively make known all the full truth of a matter
Correct The explanations should make correct predictions.
Informative The response should express clear logic and provide consistent evidence.
Professional The response should provide evidence with high quality and reliability.
Unaligned Objectives
Specific The response should refer to facts and details and avoid vague arguments.
Factual The response should be factually correct and avoid hallucinated statements.
Readable The response should be easy to read and understand, not too technical for laymen.
Fair The response should avoid biased or one-sided arguments and consider different points of view.
Repeat The response should avoid repetitive statements of one point.
Length The response should be concise and avoid redundant content.
Table 4: Text descriptions for all tested objectives.
Appendix C Experimental Details
C.1 Model Training
Details about the training process of MetaAligner and the building process of the 3 datasets are presented in Table 5. During the performance evaluation process, GPT-4 is leveraged to compare the responses on the corresponding objective. Specifically, we have the aligned test dataset: $\{q_{i},g_{i},R^{origin}_{i},R^{aligned}_{i}\}_{i=1}^{n}$, where $q$ denotes the query, $g$ denotes the ground-truth response from the dataset, $R^{origin}$ denotes the original response from the policy model, and $R^{aligned}$ denotes the aligned response from MetaAligner. We use the following prompting template and probe GPT-4 to perform judgment:
[$\mathcal{E}(q,r_{1},r_{2},\left<r\right>)$] You are a skilled evaluator of helpful AI assistants. You will be presented with one query and two different responses to this query. QUERY: {q} | RESPONSE 1: {$r_{1}$} | RESPONSE 2: {$r_{2}$}. Consider the following aspect: $r\right>$}, then return the number of the better response. If tied, return 0. You must only return 1, 2, or 0.
where $r_{1}$, $r_{2}$ are the compared response pairs, and $r\right>$ denotes the text description of the target objective. With the above information and the target objective description $r_{t}\right>$, we obtain the win rates using Algorithm 2.
C.2 Model Cards
TinyLLaMA-1.1B [36].
A compact 1.1B language model pre-trained on around 1 trillion tokens for approximately 3 epochs. Building on the architecture and tokenizer of LLaMA2, TinyLlama leverages various advances contributed by the open-source community (e.g., Flash-Attention), achieving better computational efficiency. Despite its relatively small size, TinyLlama demonstrates remarkable performance in a series of downstream tasks. It significantly outperforms existing open-source language models with comparable sizes. We use TinyLlama-1.1B-Chat-v1.0 as the base model for MetaAligner-1.1B.
LLaMA2-(Chat)-(7B, 13B, 70B) [31].
A collection of pre-trained and fine-tuned large language models (LLMs) trained and released by Meta, ranging from 7 billion to 70 billion parameters. The fine-tuned LLMs, called LLaMA2-Chat, are optimized for dialogue use cases. The models outperform other open-source models on most benchmarks. Further human evaluations prove that LLaMA2-Chat also excels in helpfulness and safety. LLaMA2 models are among the most advanced open-source foundation models. We use LLaMA2-(7B, 13B) as base models for MetaAligner-(7B, 13B), and use LLaMA2-Chat-(7B, 13B, 70B) as policy models to evaluate the alignment performances.
Vicuna-(7B, 13B, 33B) [7].
Vicuna is a family of open-source chatbots trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT. Preliminary evaluation using GPT-4 as a judge shows Vicuna-13B achieves more than 90% quality of OpenAI ChatGPT and Google Bard while outperforming other models like LLaMA and Stanford Alpaca in more than 90% of cases. We use Vicuna-(7B, 13B)-V1.5 and Vicuna-33B-V1.3 as policy models to evaluate the alignment performances.
Gemma-instruct-(2B, 7B) [30].
A family of open-source models based on Google’s Gemini models. Gemma models are pretrained on 6T tokens of text, using architectures, data, and training recipes inspired by the Gemini model family. Like Gemini, these models achieve strong generalist capabilities in text domains, alongside state-of-the-art understanding and reasoning skills at scale. Gemma-instruct models are further fine-tuned for dialogue, instruction-following, helpfulness, and safety. Gemma-instruct is developed in two sizes: a 7B version for efficient deployment and development and a 2B version for CPU and on-device applications. We select both models as policy models to evaluate the alignment performances.
MentaLLaMA-(7B, 13B, 33B) [34].
MentaLLaMA is the first open-source instruction-following LLM series for interpretable mental health analysis. Based on LLaMA2-(7B, 13B) and Vicuna-33B foundation models, MentaLLaMA is trained on the Interpretable Mental Health Instruction (IMHI) dataset with 105K instruction samples, the first multi-task and multi-source instruction-tuning dataset for interpretable mental health analysis on social media. MentaLLaMA can perform mental health analysis on social media data and generate high-quality explanations for its predictions. On evaluating sub-task performance on IMHI Benchmark (Appendix D), we introduce MentaLLaMA-(7B, 13B, 33B) models as domain-specific policy models to evaluate the alignment performances.
GPT-3.5 [22].
GPT-3.5 is an advanced, close-sourced chat-based language model developed by OpenAI. It is a sibling model to InstructGPT, which is trained to follow instructions in a prompt and provide a detailed response. The model is firstly fine-tuned with SFT with conversations in which the model played both sides—the user and an AI assistant. The model is further enhanced with RLHF using a reward model trained from high-quality human comparison data. In our experiments, we use the gpt-3.5-turbo-0125 API provided by OpenAI as a strong policy model to evaluate the alignment performances.
Claude-3 [3].
Claude-3 is among the state-of-the-art foundation models for industry benchmarks across reasoning, math, coding, multi-lingual understanding, and vision quality, developed by Anthropic. The model family includes 3 models: (1) Opus, the most capable model; (2) Sonnet, which provides a combination of skills and speed; (3) Haiku, the fastest and least expensive model. All models are multi-modal and demonstrate strong performance across benchmark evaluations. Due to the budget limits, we select claude-3-sonnet-20240229 API provided by Anthropic as a strong policy model to evaluate the alignment performances.
GPT-4 [2].
Developed by OpenAI, GPT-4 is a large-scale, multimodal foundation model that can accept image and text inputs and produce text outputs. GPT-4 marks the highest level of achievement in AI industry and exhibits human-level performance on various professional and academic benchmarks, including passing a simulated bar exam with a score around the top 10% of test takers. We leverage the strong capability of GPT-4 and use it as an oracle to evaluate the large-scale test samples. Considering the high cost of evaluating large-scale test data and our limited budget, we use the cheaper GPT-4-turbo model with the gpt-4-turbo-preview API provided by OpenAI in practice.
MetaAligner-(1.1B, 7B, 13B).
Our proposed MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment. The models are based on TinyLLaMA and LLaMA2 foundation models. We train MetaAligner models on all 3 model scales for each of the 3 benchmark datasets. Specifically, HH-RLHF-MetaAligner is trained to align the responses of a general daily AI assistant with specified objectives considering multi-turn dialogue contexts. UltraFeedback-MetaAligner is trained to align responses of another general AI assistant considering a single-turn query, but the queries include professional questions such as programming language and history, and the aligned responses are usually more complicated. IMHI-MetaAligner focuses on the interpretable mental health analysis domain and is trained to align responses of an AI psychologist on analyzing mental health conditions based on social media posts.
C.3 IMHI Annotation
We select 1,200 queries from the IMHI benchmark covering 9 mental health analysis tasks. We obtain 4 responses to each query from 4 different policy models: GPT-4-turbo [2], GPT-3.5-turbo [22], MentaLLaMA-13B [34], and LLaMA2-Chat-13B [31], with human annotations on ranking different objectives of the responses. We utilize the above policy models to generate explanations for the same query simultaneously. The annotation protocol is developed through collaborative efforts with 2 domain experts (Ph.D. students majoring in quantitative psychology) and considerations of human evaluation criteria for previous mental health analysis tasks [34, 33]. Specifically, 3 objectives are assessed: (1) Correctness: the explanations should make correct label predictions in the corresponding mental health analysis task; (2) Informativeness: the response should express clear logic and provide consistent evidence; (3) Professionalism: the response should provide evidence with high quality and reliability from the perspective of domain experts. Each aspect is divided into four standards rating from 0 to 3. Higher ratings reflect more satisfactory performance and 3 denotes approaching human performance. Each LLM-generated explanation is assigned a score by 2 domain experts for each corresponding objective, followed by the examination of 1 domain expert. All annotators are PhD students majoring in quantitative psychology.
Annotators will be given generated responses from the 4 policy models and need to score and annotate the responses from the following objectives:
Correctness. Correctness measures the trustworthiness of the classification results. Annotators should assess whether the classification result is based on facts, has misinformation, and wrong reasoning according to the given post.
  • 0: Completely unreliable information with factual hallucination (e.g. non-existent symptoms).
  • 1: Partly reliable information with wrong reasoning based on facts.
  • 2: Mostly reliable information with non-critical misinformation or wrong reasoning.
  • 3: Completely reliable information.
Informativeness. Whether the text builds from sentence to sentence to a coherent body of information and logic about mental health and supports the classification results. Annotators should assess if the generated explanation gives consistent supporting evidence to its classifications and is well-structured.
  • 0: Inconsistent with the classification results.
  • 1: Consistent with the classification results, but with poor readability and several errors.
  • 2: Consistent with the classification results. Mostly coherent and easy to read, with few minor errors.
  • 3: Consistent with the classification results. Completely fluent, coherent, and error-free.
Professionalism. Professionality measures the rationality of the generated explanations by evaluating the evidence that supports the classification results from the psychology perspective. Annotators should assess whether the explanation includes the following specified common diagnosis criteria of depression. To ensure the quality of the annotation scheme, we invite our domain experts to develop a list of common symptoms related to depression and sort these symptoms by criticality. The domain experts consult the Patient Health Questionnaire (PHQ-9) on determining the symptoms and sorting these symptoms on their knowledge.
Specifically, the following symptoms are checked (sorted by criticality):
  • Suicide ideation: Thoughts that you would be better off dead.
  • Self-harm ideation: Thoughts of hurting yourself in some way.
  • Feeling down, depressed, or hopeless.
  • Self-guilt ideation: Feeling bad about yourself — or that you are a failure or have let yourself or your family down.
  • Symptoms above are classified as with high criticality, and symptoms below are classified as with low criticality.
  • Feeling tired or having little energy. Little interest or pleasure in doing things.
  • Poor appetite or overeating.
  • Trouble falling or staying asleep, or sleeping too much.
  • Trouble concentrating on things, such as reading the newspaper or watching television.
  • Moving or speaking so slowly that other people could have noticed. Or the opposite — being so fidgety or restless that you have been moving around a lot more than usual
  • Uncontrollable sexual desire or sexual frigidity.
  • Other symptoms.
Based on the above symptoms, the annotators score the professionality of each explanation with the following criteria:
  • 0: The explanation provides no supportive evidence or symptoms with high criticality are missing in the explanation.
  • 1: The explanation provides a few supportive evidence, while some symptoms with higher criticality (than provided evidence) are missing.
  • 2: The explanation provides several supportive evidence, while some symptoms with lower criticality (than provided evidence) are missing.
  • 3: The explanation provides all related supportive evidence in the post.
Algorithm 2 GPT-4 win-rate computation
1:The aligned test dataset: $\{q_{i},g_{i},R^{origin}_{i},R^{aligned}_{i}\}_{i=1}^{n}$; Text description for target objective: $r_{t}\right>$; Prompting template: $\mathcal{E}(q,r_{1},r_{2},\left<r\right>)$
2:Win rate of the aligned responses $\omega$.
3:$W_{origin}\leftarrow\varnothing;W_{aligned}\leftarrow\varnothing$triangleright Initialize the judgement set W.
4:$P_{origin}\leftarrow\varnothing;P_{aligned}\leftarrow\varnothing$triangleright Initialize the set P to record the position of the responses.
5:$win_{origin}=0;win_{aligned}=0$triangleright Initialize the counter for wining samples.
6:for $i\in\{1,...,n\}$ do
7:     $r_{1}^{origin},r_{2}^{origin},p^{origin}_{i}=\textit{random\_shuffle}(g_{i},R^{origin}_{i})$triangleright Random shuffle the origin and ground-truth response. $p^{origin}_{i}$ denotes the position of $R^{origin}_{i}$.
8:     $P_{origin}\leftarrow\mathcal{E}(q_{i},r_{1}^{origin},r_{2}^{origin},\left<r_{t}\right>)$triangleright Prompt for comparing origin and ground-truth response.
9:     $J^{origin}_{i}\leftarrow\textit{Call-GPT-4}(P_{origin})$triangleright Call GPT-4 API to perform judgement.
10:     $W_{origin}\leftarrow W_{origin}\cup\{J^{origin}_{i}\}$
11:     $P_{origin}\leftarrow P_{origin}\cup\{p^{origin}_{i}\}$
12:     if $J^{origin}_{i}=p^{origin}_{i}$ then
13:         $win_{origin}=win_{origin}+1$
14:     end if
15:     $r_{1}^{aligned},r_{2}^{aligned},p^{aligned}_{i}=\textit{random\_shuffle}(g_{i},R^{aligned}_{i})$$\triangleright$ Similar actions for aligned response.
16:     $P_{aligned}\leftarrow\mathcal{E}(q_{i},r_{1}^{aligned},r_{2}^{aligned},\left<r_{t}\right>)$
17:     $J_{aligned}\leftarrow\textit{Call-GPT-4}(P_{aligned})$
18:     $W_{aligned}\leftarrow W_{aligned}\cup\{J^{aligned}_{i}\}$
19:     $P_{aligned}\leftarrow P_{aligned}\cup\{p^{aligned}_{i}\}$
20:     if $J^{aligned}_{i}=p^{aligned}_{i}$ then
21:         $win_{aligned}=win_{aligned}+1$
22:     end if
23:end for
24:$\omega_{origin}=\frac{win_{origin}}{\textit{len}(W_{origin})}$triangleright Calculate win rates for original responses over ground-truth responses.
25:$\omega_{aligned}=\frac{win_{aligned}}{\textit{len}(W_{aligned})}$triangleright Calculate win rates for aligned responses over ground-truth responses.
26:$\omega=\omega_{aligned}-\omega_{origin}$
Training Information
Base Library Huggingface Transformers
Fine-tuning Platform FastChat
GPU Hardware 4$\times$ NVIDIA Tesla A100 80GB GPUs
CPU Hardware 8$\times$ Intel(R) Xeon(R) Gold 6342 CPU cores per GPU
Hardware Speedup Flash Attention 2 [9]
Quantization for training BF16
Fine-tuning Strategy Full fine-tuning
Optimizer Adam
Training Epochs 2
Batch sizes HH-RLHF: 512 / UltraFeedback: 512 / IMHI: 128
Max token for training MetaAligner-(1.1B, 7B, 13B): 2048/4096/4096
Learning rate 1e-5
Warm-up ratio 0.05
Base Model-1.1B TinyLLaMA-1.1B
Base Model-7B/13B LLaMA2-Chat-(7B, 13B)
Dataset Information
Dataset Name HH-RLHF
Licence MIT
Train/Val/Test ($\mathcal{D}_{p}$) 262,719/15,000/15,000
Train/Val ($\mathcal{D}_{e}$) 16,502/1,797
Harmless preference source Ray2333/gpt2-large-harmless-reward_model
License MIT
Helpful preference source Ray2333/gpt2-large-helpful-reward_model
License MIT
Humor preference source mohameddhiab/humor-no-humor
License Apache-2.0
Test evaluator GPT-4
Dataset Name UltraFeedback
Licence MIT
Train/Val/Test ($\mathcal{D}_{p}$) 252,934/15,000/15,000
Train/Val ($\mathcal{D}_{e}$) 82,023/5,000
Instruction_following preference source GPT-4
Honest preference source GPT-4
Truthful preference source GPT-4
Helpful preference source GPT-4
Test evaluator GPT-4
Dataset Name IMHI
Licence MIT
Train/Val/Test ($\mathcal{D}_{p}$) 5,304/1,051/2,400
Train/Val ($\mathcal{D}_{e}$) 3,374/689
Instruction_following preference source Human annotation
Correct preference source Human annotation
Informative preference source Human annotation
Professional preference source Human annotation
Test evaluator GPT-4
Policy Models
LLaMA2-Chat-(7B, 13B, 70B) https://huggingface.co/meta-llama
Gemma-instruct-(2B, 7B) https://huggingface.co/google
Vicuna-(7B, 13B, 33B) https://huggingface.co/lmsys
GPT-3.5 https://openai.com/blog/chatgpt
Claude-3-Sonnet https://www.anthropic.com/news/claude-3-family
Table 5: Details for MetaAligner training and datasets. ‘preference source’ denotes how the preference annotations were obtained.
Appendix D Sub-task Performance on IMHI Benchmark
Data Task Source Annotation Labels/Aspects
DR depression detection Reddit weak supervision Yes, No
Dreaddit stress detection Reddit human annotation Yes, No
SWMH mental disorders detection Reddit weak supervision Suicide, Anxiety, Bipolar disorder, Depression, None
T-SID mental disorders detection Twitter weak supervision None, Suicide, Depression, PTSD
SAD stress cause detection SMS human annotation
School, Finance, Family, Social Relation,
Work, Health, Emotion, Decision, Others
CAMS depression/suicide cause detection Reddit human annotation
Bias, Jobs, Medication, Relationship,
Alienation, None
loneliness loneliness detection Reddit human annotation Yes, No
MultiWD Wellness dimensions detection Reddit human annotation
Spiritual, Physical, Intellectual, Social,
Vocational, Emotional
IRF interpersonal risk factors detection Reddit human annotation Thwarted Belongingness, Perceived Burdensomeness
Table 6: Details about the 9 sub-tasks in the IMHI dataset. "Annotation" denotes the reliability of the annotations in the raw data.
MetaAligner Policy Model CAMS DR Dreaddit IRF loneliness MultiWD SAD SWMH T-SID Overall
1.1B LLaMA2-Chat-7B -3.4% +28.0% +12.67% +45.67% +38.0% +34.67% +23.33% +33.0% +31.33% +27.04%
LLaMA2-Chat-13B -25.0% +27.67% +22.67% +52.33% +32.0% +31.33% +23.33% +23.67% +36.67% +29.7%
LLaMA2-Chat-70B +5.33% +35.0% +19.0% +46.33% +42.0% -3.0% +8.0% +0.33% +3.0% +21.18%
Gemma-instruct-2B +32.0% +4.33% +37.0% +40.0% +34.0% +61.0% +45.0% +52.0% +27.33% +38.77%
Gemma-instruct-7B +16.33% +51.33% +42.67% +44.67% +51.0% +55.33% +49.33% +40.0% +53.67% +44.92%
MentaLLaMA-7B +13.33% +22.67% +23.33% +47.33% +39.67% +39.33% +33.33% +30.67% +41.0% +32.29%
MentaLLaMA-13B +21.34% +31.0% +39.0% +29.67% +42.33% +47.66% +20.67% +34.33% +33.0% +36.03%
MentaLLaMA-33B -5.44% -0.33% -5.33% +30.67% +3.0% +14.0% +24.33% +7.0% +2.33% +6.66%
Vicuna-7B +6.0% -9.33% +24.0% +70.67% +17.67% +33.33% +17.33% +22.0% +42.0% +24.85%
Vicuna-13B +6.0% -9.0% +13.67% +71.67% +20.33% +32.67% +21.0% +2.0% +16.67% +24.97%
Vicuna-33B +29.33% +2.0% -1.0% +41.0% +10.0% +9.67% +1.33% -9.67% +1.33% +9.33%
7B LLaMA2-Chat-7B +7.67% +47.0% +13.0% +33.0% +38.33% +31.0% +17.0% +38.67% +49.33% +30.55%
LLaMA2-Chat-13B -12.34% +51.67% +18.0% +37.67% +43.33% +29.33% +22.33% +18.0% +38.34% +27.37%
LLaMA2-Chat-70B +15.66% +40.53% +9.0% +28.33% +25.34% +6.67% +31.33% +10.33% +10.66% +20.67%
Gemma-instruct-2B +45.0% +8.33% +30.0% +57.0% +39.0% +48.66% +57.0% +52.0% +36.33% +41.11%
Gemma-instruct-7B +32.0% +42.67% +20.67% +45.0% +36.33% +43.0% +46.0% +40.67% +20.7% +38.0%
MentaLLaMA-7B +20.66% +50.67% +19.66% +35.66% +30.0% +27.0% +40.33% +36.34% +41.0% +35.48%
MentaLLaMA-13B +25.0% +45.66% +43.66% +34.34% +36.33% +30.33% +48.0% +42.33% +36.0% +37.97%
MentaLLaMA-33B -5.33% +2.66% -9.66% +2.34% +21.33% -4.67% +20.0% +8.67% +3.33% +4.22%
Vicuna-7B +22.33% +7.67% +4.0% +40.0% +20.0% +14.0% +6.0% +7.67% +15.0% +15.19%
Vicuna-13B +1.0% +48.0% +8.0% +67.0% +33.0% +12.0% +15.0% +23.0% +22.0% +32.33%
Vicuna-33B -2.0% +54.0% +10.0% +62.0% +37.0% +29.0% +10.0% +15.0% +12.0% +25.22%
13B LLaMA2-Chat-7B +27.33% +45.0% +23.33% +62.67% +54.33% +59.34% +52.0% +55.33% +70.33% +49.96%
LLaMA2-Chat-13B +5.34% +47.0% +22.67% +65.33% +56.33% +52.33% +31.33% +32.0% +56.67% +35.55%
LLaMA2-Chat-70B +28.0% +59.0% +23.33% +57.33% +53.0% +7.0% +27.33% +23.0% +23.67% +31.18%
Gemma-instruct-2B +52.66% +19.66% +41.33% +60.67% +50.67% +79.0% +53.67% +63.33% +45% +51.74%
Gemma-instruct-7B +35.33% +52.0% +39.0% +55.34% +50.0% +64.0% +56.33% +49.33% +61.0% +50.36%
MentaLLaMA-7B +40.67% +38.34% +29.33% +65.67% +38.34% +57.0% +46.66% +49.0% +61.67% +47.4%
MentaLLaMA-13B +34.0% +27.0% +39.0% +54.33% +35.33% +49.33% +41.67% +46.67% +45.66% +43.62%
MentaLLaMA-33B +2.67% -8.0% +2.67% +30.67% -8.0% +23.67% +29.33% +28.67% +20.0% +11.2%
Vicuna-7B +23.33% +17.67% +8.0% +63.67% +35.67% +36.33% +12.33% +18.0% +37.0% +28.07%
Vicuna-13B +15.0% -7.33% +2.0% +67.67% +24.33% +40.67% +19.0% +4.0% +17.67% +20.29%
Vicuna-33B -6.0% +54.0% +36.0% +79.0% +38.0% +29.0% +24.0% +21.0% +24.0% +33.22%
Table 7: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models. The GPT-4 judge considers 3 objectives: "Correct", "Informative", and "Professional". The figures show the advantage of aligned outputs over the policy model outputs on win rate. Best values for each MetaAligner model are highlighted in bold.
We stack MetaAligner on different policy models to perform alignment on all 3 objectives: "Correct", "Informative", and "Professional". We include MentaLLaMA-(7B, 13B, 33B) [34], the first open-source instruction-following LLM series for interpretable mental health analysis into the policy models. Details about the 9 sub-tasks are provided in Table 6. The overall performance of MetaAligner on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 7.
According to the results, the MetaAligner models achieve substantial improvement in overall performance on all 11 policy models, with an average of 26.89% advantage on win rates for MetaAligner-1.1B, 28.01% for MetaAligner-7B, and 36.6% for MetaAligner-13B. These results show the general effectiveness of one MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment. MetaAligner also greatly improves performance on each sub-task. For example, MetaAligner-7B outperforms the unaligned outputs by over 25% on 7 sub-tasks. These results indicate that MetaAligner alignment can be effectively adapted to tasks that require different knowledge and response formats.
From the policy model scale perspective, MetaAligner provides successful alignments to models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself. In the extreme case, MetaAligner-1.1B advances the win-rate of LLaMA2-Chat-70B outputs by 21.18%, a policy model with 63$\times$ more parameters. These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow. Besides the general-domain foundation models, MetaAligner also improves the performance by an average of 28.32% on MentaLLaMA models, which are fine-tuned on mental health analysis tasks. These results show that MetaAligner can make reasonable corrections on weak responses while maintaining their expertise from domain-specific policy models.
Appendix E Objective-wise Alignment
E.1 Experimental Settings
We randomly sample 1,200 queries from the UltraFeedback test set and probe the target policy models to provide responses to all queries, which are regarded as unaligned responses. Then MetaAligner is used to align these responses under different objectives, including aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and another full alignment on all 4 objectives. During evaluation, GPT-4 is leveraged as a reward model to score these responses considering different objectives. Specifically, the following prompt is developed to obtain the scores:
[$\mathcal{S}(q,a,\left<r\right>)$] You are a skilled evaluator of helpful AI assistants. You will be presented with one query and a response to this query. QUERY: {q} | RESPONSE: {a} Assign a score ranging from 0 to 10 to this response considering the following aspect: $r\right>$}. The rating improves as the score rises, where 0 denotes inferior performance and 10 denotes approaching-human performance. You must only return the assigned score.
where $q$ and $a$ denotes the target query and response, and $r\in${Instruction following, Helpful, Honest, Truthful} is the target objective for evaluation.
E.2 Experimental Results
More experimental results are presented in Figure 4. The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5 outputs from the UltraFeedback test set. According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5. Unlike full alignment which shows weaker performance as the capability of the policy model increases, objective-wise alignment provides stable improvement for different policy models. For example, aligning "Instruction following" leads to over 20% of approaching-human responses for all policy models, and aligning "Honesty" leads to over 30% of approaching-human responses for all policy models. The reason is that single-objective alignment doesn’t involve the complex interactions between multiple objectives, which facilitates the full realization of MetaAligner’s potential to improve the response on the corresponding objective.
Figure 4: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
Appendix F Implementation Details of Baseline Models
F.1 SFT-based Methods
Existing SFT-based methods [11, 35] for multi-objective preference alignment explicitly incorporate reward values into the query via prompt engineering, which includes a text marker for each alignment objective and their corresponding value numbers in the current response. The model is trained to predict the response given the enhanced query, which enables it to learn a mapping between the reward value and its reflection in the response. During inference, we achieve alignment by assigning a higher reward value to the target objectives. Specifically, we define the following prompting template for HH-RLHF dataset:
[$\mathcal{F}(q,r_{1},r_{2},r_{3}$] <Harmlessness>: {$r_{1}$}; <Helpfulness>: {$r_{2}$}; <Humour>: {$r_{3}$} | {q}
where $q$ denotes the query, $r_{1}$, $r_{2}$, $r_{3}$ denote the corresponding reward values for the current response, ranging from 1 to 5. We define the following prompting template for UltraFeedback dataset:
[$\mathcal{F}(q,r_{1},r_{2},r_{3},r_{4}$] <instruction_following>: {$r_{1}$}; <honesty>: {$r_{2}$}; <honesty>: {$r_{3}$}; <helpfulness>: {$r_{4}$} | {q}
where $q$ denotes the query, $r_{1}$, $r_{2}$, $r_{3}$, $r_{4}$ denote the corresponding reward values for the current response, obtained from existing reward models.
During inference, we aim to simultaneously align all objectives with 1 model to enable fair comparisons to MetaAligner. For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during inference: $r_{1}=r_{2}=r_{3}=r_{4}=5$. For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models. Therefore, we set each objective value to its maximum in the training dataset to enable higher alignment. The values are set to: $r_{1}=4.19$, $r_{2}=3.03$, $r_{1}=1.58$.
In calculating the GPU hours, we sum the training hours for fine-tuning the LLaMA-Chat-7B policy model for HH-RLHF and UltraFeedback datasets.
F.2 MODPO
We implement MODPO by implementing the Controllable Direct Preference Optimization (CDPO) [11] algorithm based on its paper introduction. We bypass the controllable preference SFT stage by utilizing the trained model from SFT-based methods as the foundation model. In the CDPO stage, each query $q$ is accompanied by two responses $c_{1}$ and $c_{2}$, where each response is constructed into contrastive pairs. In HH-RLHF, we have the following prompting template:
[$\mathcal{F}(q,c_{1},r_{1}^{1},r_{1}^{2},r_{1}^{3}$] <Harmlessness>: {$r_{1}^{1}$}; <Helpfulness>: {$r_{1}^{2}$}; <Humour>: {$r_{1}^{3}$} | {q} | {$c_{1}$}
[$\mathcal{F}(q,c_{2},r_{2}^{1},r_{2}^{2},r_{2}^{3}$] <Harmlessness>: {$r_{2}^{1}$}; <Helpfulness>: {$r_{2}^{2}$}; <Humour>: {$r_{2}^{3}$} | {q} | {$c_{2}$}
where $r_{i}^{j}$ denotes the reward value for the $i$-th response on the $j$-th objective. For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models. Therefore, we set each preference value to its maximum in the training dataset to enable higher alignment. The values are set to: $r^{1}=4.19$, $r^{2}=3.03$, $r^{3}=1.58$. Similarly, on UltraFeedback we have:
[$\mathcal{F}(q,c_{1},r_{1}^{1},r_{1}^{2},r_{1}^{3},r_{1}^{4}$] <instruction_following>: {$r_{1}^{1}$}; <honesty>: {$r_{1}^{2}$}; <honesty>: {$r_{1}^{3}$}; <helpfulness>: {$r_{1}^{4}$} | {q} | {$c_{1}$}
[$\mathcal{F}(q,c_{2},r_{2}^{1},r_{2}^{2},r_{2}^{3},r_{2}^{4}$] <instruction_following>: {$r_{2}^{1}$}; <honesty>: {$r_{2}^{2}$}; <honesty>: {$r_{2}^{3}$}; <helpfulness>: {$r_{2}^{4}$} | {q} | {$c_{2}$}
For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during the DPO training process: $r^{1}=r^{2}=r^{3}=r^{4}=5$. In determining the preference scores, we use the CDPO learning goal to transform the task into a conditional multi-objective optimization problem. Specifically, the reward value $R_{i}$ for response $c_{i}$ is calculated as follows:
$R_{i}=\sum_{i=1}^{m}\omega_{i}g_{i}$ (9)
where $\omega_{i}$ represents the weight of the $i$-th objective and $g_{i}$ is calculated as follows:
$\small g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (10)
where $\lambda_{i}$ represents the weight of the controlled objective and $p_{j}$ is a pre-defined preference value for the $j$-th objective. In our implementation, we set $\omega_{i}=\lambda_{i}=1$. We aim to simultaneously align all objectives with 1 model to enable fair comparisons to MetaAligner. For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during inference: $p_{1}=p_{2}=p_{3}=p_{4}=5$. For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models. Therefore, we set each preference value to its maximum in the training dataset to enable higher alignment. The values are set to: $p_{1}=4.19$, $p_{2}=3.03$, $p_{3}=1.58$. After calculation, the response with a higher reward value $R_{i}$ is used as the chosen response, and the other response is used as the rejected response for MODPO training. Specifically, the model is trained via the following loss function:
$\mathcal{L}_{\mathrm{CDPO}}={\tiny-\mathbb{E}_{\left(x,c,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid c,x)}{\pi_{\text{ref }}(y_{w}\mid c,x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid c,x)}{\pi_{\text{ref }}(y_{l}\mid c,x)}\right)\right]}$ (11)
where $x$ denotes the query, $y_{w},y_{l}$ denote the chosen and rejected prompts, $c$ denotes the corresponding condition, $\pi_{\theta}$ and $\pi_{ref}$ denote the target policy model and the reference model. For implementation, we build our MODPO code based on the OpenRLHF library.
In calculating the GPU hours, we include the training hours for tuning the SFT-based policy model using the CDPO algorithm for HH-RLHF and UltraFeedback datasets. We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for fair comparisons.
F.3 MORLHF
We use the linear scalarization [27, 20] realization of MORLHF with the KL-divergence regularization. Specifically, the model is trained via the following objective function:
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{ref}(y|q)}\right]$ (12)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector, $q$, $y$ denote the query and the response, R denotes the reward functions for the target objectives. In implementing the MORLHF algorithm, we first train a reward model for each objective based on random samples from 50% of the training data. Following the reward models we used in building the HH-RLHF dataset, we select the GPT2-large as foundation models for the reward model, and optimize the reward models using the following pair-wise loss functions:
$\mathcal{L}_{rm}=-log(\sigma(R_{c}-R_{r}-margin))$ (13)
where $\sigma$ denotes the Sigmoid function, $R_{c}$ and $R_{r}$ denote the reward output of the chosen response and the rejected response, and $margin$ denotes the margin loss for the corresponding response pairs when multiple responses are ranked, such as in UltraFeedback. Secondly, we fine-tune the LLaMA2-Chat-7B policy model with the highest-ranked responses from the HH-RLHF and UltraFeedback datasets to obtain sub-optimal starting points for RLHF. The SFT training process is formalized as follows:
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q)\right]$ (14)
where $q$ and $y$ denote the query and its corresponding highest-ranked response. Thirdly, following most works in RLHF, we leverage the PPO algorithm [26] to enable parameterized optimization of the policy model. In linear scalarization, we set $\omega_{1}=\omega_{2}=...=\omega_{N}=\frac{1}{N}$. For implementation, we build our MORLHF code based on the OpenRLHF library.
In calculating the GPU hours, we include the training hours for all reward models in HH-RLHF and UltraFeedback datasets, with a sum of seven reward model training processes. We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for reaching the sub-optimal starting points. Finally, the PPO training hours for HH-RLHF and UltraFeedback are included in the GPU hours.